By the addition theorem of probability:
$$P(A \cup B) = P(A) + P(B) - P(A \cap B) = 0.25 + 0.50 - 0.14 = 0.61$$The event that 'neither $A$ nor $B$ occurs' is the complement of $(A \cup B)$:
$$P(\text{neither } A \text{ nor } B) = P((A \cup B)^c) = 1 - P(A \cup B) = 1 - 0.61 = 0.39$$प्रायिकता के योग प्रमेय से:
$$P(A \cup B) = P(A) + P(B) - P(A \cap B) = 0.25 + 0.50 - 0.14 = 0.61$$'न तो $A$ और न ही $B$ के घटित होने' की प्रायिकता $(A \cup B)$ की पूरक घटना है:
$$P((A \cup B)^c) = 1 - P(A \cup B) = 1 - 0.61 = 0.39$$